OpenStudy (anonymous):

Solve the equation: l-4+3xl=14

OpenStudy (anonymous):

4+3x=14 4+3x=-14 3x=10 3x=-18 x=10/3 x=-18/3

OpenStudy (anonymous):

|-4+3x|=14 Remove the absolute value term. This creates a \ on the right-hand side of the equation because |x|=\x. -4+3x=\(14) Set up the + portion of the \ solution. -4+3x=14 Move all terms not containing x to the right-hand side of the equation. 3x=18 Divide each term in the equation by 3. x=6 Set up the - portion of the \ solution. -4+3x=-(14) Multiply -1 by the 14 inside the parentheses. 3x-4=-14 Since -4 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4 to both sides. 3x=4-14 Subtract 14 from 4 to get -10. 3x=-10 Divide each term in the equation by 3. (3x)/(3)=-(10)/(3) Simplify the left-hand side of the equation by canceling the common factors. x=-(10)/(3) The solution to the equation includes both the positive and negative portions of the solution. x=6,-(10)/(3)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

any more question u want to ask

OpenStudy (anonymous):

So you basically just put a negative sign on the opposite side of the equals sign for the second answer? Or is there more to it?